An earth satellite moves in a circular orbit at a speed of 5,800 m/s. What is its orbital period in minutes?
Answer #1

orbital distance

v^2 = GM / R

5800^2 = 6.67* 10^-11* 5.98* 10^24 / R

R = 1.1857* 10^7 m

now

T = 2 pi R / v

T = 213.97 mins

=====

Comment before rate in case any doubt, will reply for sure.. goodluck